# On the Mode-Seeking Properties of Langevin Dynamics

### v2

@@ 6 Numerical Results
In this section, we empirically evaluated the mode-seeking tendencies of vanilla and chained Langevin dynamics. We performed numerical experiments on synthetic Gaussian mixture models and real image datasets including MNIST [41] and Fashion-MNIST [42]. Details on the experiment setup are deferred to Appendix D.
Synthetic Gaussian mixture model: We consider the data distribution $P$ as a mixture of three Gaussian components in dimension $d=100$, where mode 0 defined as $P^{(0)}={\mathcal{N}}(\bm{0}_{d},3{\bm{I}}_{d})$ is the in-between mode with high variance, and mode 1 and mode 2 are respectively defined as $P^{(1)}={\mathcal{N}}(\bm{1}_{d},{\bm{I}}_{d})$ and $P^{(2)}={\mathcal{N}}(-\bm{1}_{d},{\bm{I}}_{d})$. The frequencies of the three modes are 0.2, 0.4 and 0.4, i.e.,
@@
In the synthetic experiments, we give the samplers access to the true score function calculated from the target distribution. As shown in Figure 2, vanilla Langevin dynamics (VLD) cannot find mode 1 or 2 within $10^{6}$ iterations if the sample is initialized in mode 0, while chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=10$ can find the other two modes in 1000 steps and correctly recover their frequencies as gradually increasing the number of iterations. When the sample is initialized in mode 1, as shown in Figure 5 in Appendix D.1, VLD is also likely to be trapped by the high-variance mode 0 and cannot find mode 2, while Chained-VLD is capable of finding all modes. Additional experiments on samples initialized in mode 2 are presented in Appendix D.1, which also verify the mode-seeking tendencies of vanilla Langevin dynamics. We also investigated the effect of different choices of patch size $Q$ on the performance of Chained-LD. As shown in Figures 7, 8, and 9 in Appendix D.1, the convergence of Chained-LD are insensitive to moderate values of constant $Q\in\{1,4,10\}$; for large $Q=20$, it takes more steps to find the other modes; while for overly large $Q=50$, Chained-LD has mode-seeking tendencies similar to LD.
@@
Image datasets: We also perform experiments on generating samples from image datasets by chained annealed Langevin dynamics (Chained-ALD). We construct the distribution as a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. randomly flipping an image (white background and black digits/objects) with probability 0.5.
Since the target distribution of image datasets is unknown, following from [4], we train an estimator to approximate the score function from training samples. More details are deferred to Appendix D.2. We use Recurrent Neural Network (RNN) architectures to estimate the perturbed conditional score function $\nabla_{{\mathbf{x}}^{(q)}}\log P_{{\sigma}_{t}}({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$ for Chained-ALD. We note that for a sequence of inputs, the output of RNN from the previous step is fed as input to the current step. Therefore, in the scenario of chained Langevin dynamics, the hidden state of RNN contains information about the previous patches and allows the network to estimate the conditional score function of the next patch. More implementation details are deferred to Appendix D.3.
@@
We numerically compare the performance of annealed Langevin dynamics (ALD) and Chained-ALD with different noise levels. The experimental results are shown in Figures 3 and 4. For ALD with bounded noise levels (i.e., the maximum noise $\sigma_{\max}=1$), we observe that it tends to generate the samples from the same mode as initialization, aligning with our theoretical analysis in Theorem 4 in Appendix B. Then, if we apply significantly larger noise levels (i.e., the maximum noise $\sigma_{\max}=50$ as suggested by Technique 1 in [26]), ALD could generate samples from both modes. On the other hand, Chained-ALD, even with bounded noise levels (i.e., $\sigma_{\max}=1$), is capable of finding both modes. Further experiments are deferred to Appendix D.3.
@@ 7 Conclusion
In this work, we theoretically and numerically studied the mode-seeking properties of Langevin dynamics sampling methods under a multi-modal distribution. We characterized Gaussian and sub-Gaussian mixture models under which vanilla Langevin dynamics are unlikely to find all the components within a sub-exponential number of iterations. To reduce the mode-seeking tendency of vanilla Langevin dynamics, we proposed Chained Langevin Dynamics (Chained-LD) and analyzed its convergence behavior. Studying the connections between Chained-LD and denoising diffusion models will be an interesting topic for future exploration. Our RNN-based implementation of Chained-LD is currently limited to image data generation tasks. An interesting future direction is to extend the application of Chained-LD to other domains such as audio and text data. Another future direction could be to study the convergence of Chained-LD under an imperfect score estimation.
Appendix A Theoretical Analysis on the Mode-Seeking Tendency of Langevin Dynamics

### v3

@@ 6 Numerical Results
In this section, we empirically validated our theoretical findings of vanilla and chained Langevin dynamics. We performed numerical experiments on synthetic Gaussian mixture models and real image datasets including MNIST (lecun1998mnist, ) and Fashion-MNIST (xiao2017fashion, ). Details on the experiment setup are deferred to Appendix D.
Synthetic Gaussian mixture model: We consider the data distribution $P$ as symmetric 3-Gaussian model with $w=0.2$, $\nu=\sqrt{3}$, and ${\bm{\mu}}=\bm{1}_{d}$, i.e.,
$P=0.2P^{(0)}+0.4P^{(1)}+0.4P^{(2)}=0.2{\mathcal{N}}(\bm{0}_{d},3{\bm{I}}_{d})+0.4{\mathcal{N}}(\bm{1}_{d},{\bm{I}}_{d})+0.4{\mathcal{N}}(-\bm{1}_{d},{\bm{I}}_{d}).$ (5)
@@
In the synthetic experiments, we give the samplers access to the true score function calculated from the target distribution. As shown in Figure 3, vanilla Langevin dynamics (VLD) cannot find mode 1 or 2 within $10^{6}$ iterations if the sample is initialized in mode 0, while chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=10$ can find the other two modes in 1000 steps and correctly recover their frequencies as gradually increasing the number of iterations. When the sample is initialized in mode 1, as shown in Figure 5 in Appendix D.1, VLD is also likely to be trapped by the high-variance mode 0 and cannot find mode 2, while Chained-VLD is capable of finding all modes. Additional experiments on samples initialized in mode 2 are presented in Appendix D.1, which also verify the convergence hardness of vanilla Langevin dynamics. We also investigated the effect of different choices of patch size $Q$ on the performance of Chained-LD. As shown in Figures 7, 8, and 9 in Appendix D.1, the convergence of Chained-LD are insensitive to moderate values of constant $Q\in\{1,4,10\}$; for large $Q=20$, it takes more steps to find the other modes; while for overly large $Q=50$, Chained-LD has convergence hardness similar to LD.
@@
Applications of Chained-LD in generative modeling: We also test the application of Chained-LD as a sampling methodology in generative modeling. We consider a mixture distribution of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. randomly flipping an image (white background and black digits/objects) with probability 0.5. Following from (song2019generative, ), we train an estimator to approximate the score function from training samples, and apply Chained-LD using the estimated score function. More implementation details are deferred to Appendix D.2.
@@
We numerically validate our theoretical findings of annealed Langevin dynamics (ALD) and Chained-ALD. As shown in Figures 4, ALD with bounded noise levels (i.e., the maximum noise $\sigma_{\max}=1$) tends to sample from the same mode as initialization, aligning with our theoretical analysis in Theorem 3. Then, if we apply larger noise levels (i.e., the maximum noise $\sigma_{\max}=50$ as suggested by Technique 1 in (song2020improved, )), ALD could generate samples from both modes. On the other hand, Chained-ALD, even with bounded noise levels (i.e., $\sigma_{\max}=1$), is capable of finding both modes. Further experiments are deferred to Appendix D.3.
@@ 7 Conclusion
In this work, we theoretically and numerically studied the hardness of Langevin dynamics sampling methods under a multi-modal distribution. We characterized Gaussian and sub-Gaussian mixture models under which Langevin dynamics are unlikely to find all the components within a sub-exponential number of iterations. To reduce the exponential iteration complexity of Langevin dynamics, we proposed Chained Langevin Dynamics (Chained-LD), as a complementary solution to Annealed-LD in (song2019generative, ) and analyzed its convergence behavior. Further investigation on the applications of Chained-LD in generative models will be an interesting topic for future exploration. Another future direction could be to study the convergence of Chained-LD under an imperfect score estimation.
Appendix A Iteration Complexity of Langevin Dynamics in Gaussian Mixture Models